Membership in an elite oraganization requires a test score in the upper 30% range. If the mean is 115 and the standard deviation is 12,find the lowest acceptable score that would enable a candidate to apply for membership. Assume the variable is normally distributed.
Answer by stanbon(75887)

Membership in an elite oraganization requires a test score in the upper 30% range. If the mean is 115 and the standard deviation is 12,find the lowest acceptable score that would enable a candidate to apply for membership. Assume the variable is normally distributed.
====
Find the z-score with a left-tail 0 70%.
invNorm(0.70) = 0.5244
----
Find the corresponding test score using x = z*s + u
x = 0.5244*12 + 115 = 121.29
======================
Cheers,
Stan H.